# AutoCodeRover: Autonomous Program Improvement

### v1

@@ 6.2. RQ2: Effect of SBFL
We present a case study on one of the tasks uniquely resolved by ACR-val-sbfl. Figure 7 shows the issue description of django-13964666https://code.djangoproject.com/ticket/32332 and how AutoCodeRover attempts to retrieve code context with and without SBFL. This issue reported a bug when saving django models to a database. A simplified version of the issue description is shown in the first part of Figure 7, in which important parts are highlighted. In this issue, some “hints” (highlighted in red) mentioned are actually distracting factors for a context retrieval agent. For example, the Product and Order classes describe how the bug can be reproduced, and are not classes that cause the bug. Retrieving code context from these two classes (as shown in Figure 7 part 2 “WITHOUT SBFL”) does not yield useful results for resolving the issue. In all AutoCodeRover variants without SBFL, the context retrieval agent invokes similar APIs as presented in Figure 7, and thus could not gather sufficient code context for patch generation.
@@
method is actually where the developer chose to fix this bug [^7] https://github.com/django/django/pull/13964/files). Moreover, we observe that the agent does not solely rely on the SBFL results to make API invocations. Instead of searching for methods ranked as top-1 in the SBFL results, the agent searched for the 3rd and 5th ranked methods. These methods are more related to some other hints mentioned in the issue description (highlighted in yellow), and the LLM agent is able to exploit this correlation between the natural language descriptions and the SBFL analysis results. With the correct context collected, AutoCodeRover is then able to draft a patch that resolves this issue. This investigation suggests that an execution-based analysis can complement the agent workflow by revealing more information not included in the issue description, thus improving its performance.
@@ 6.3. RQ3: Challenges on real-life tasks
In this research question, we analyze the task instances in SWE-bench lite that AutoCodeRover failed to resolve, and provide a taxonomy of the issue characteristics to highlight the practical challenges in achieving fully automated software improvement. Our taxonomy consists of challenges in the fault localization stage and patch generation stage. Specifically, for each task instance, we analyze the best run in the three repetitions (i.e. ACR-{1-3}), and classify each task instance into one of the following:
@@
Figure 8 shows the distribution of the 300 tasks instances in SWE-bench lite. AutoCodeRover resolves 22.3% of the issues (“Success”), as mentioned in Section 6.1. The fail-to-resolve cases are included in the remaining four categories. In 31.7% of the tasks, AutoCodeRover correctly decided on all patch locations (at the method-level), but did not produce a correct patch (“Wrong patch”). More fine-grained intra-procedural analysis and specification inference techniques can play a significant role in improving these cases, by providing the patch generation agent with more method-level repair guidance.
In the other three categories, the fault localization could not pinpoint all the locations to be modified. In 20.7% of the tasks, a patch is generated in the correct file, but at wrong methods / classes in the file (“Wrong location in correct file”). In some of these runs, the developer patch modifies multiple methods, but the generated patch did not modify all of them. In the other categories, a patch could not be generated at the correct file - in 18.7% of the tasks a patch is generated in wrong files, and in 6.7% of the tasks there is no applicable patch (“Wrong file” and “No patch”). We manually inspected some tasks in these two categories, and observed that their issue description mentions few methods / classes / files in the codebase. Instead, some of them contain short examples to reproduce the issue. For these tasks, one possibility is to generate a comprehensive test-suite based on the issue description, and then use execution information of the test-suite (e.g. SBFL) to reveal suspicious program locations. On the other hand, some other tasks do not contain reproducible examples and only consist of natural language descriptions. This is where some level of human involvement might be helpful - e.g. human developers can provide some initial directions and hints for the AI-based tools.
@@ 8. Threats to Validity
Despite the high efficacy AutoCodeRover achieved in SWE-bench lite, we report a few potential threats to our approach and experiments and discuss how we addressed them. First, LLM is known for its randomness to generate different results in different runs, which may threaten the validity of AutoCodeRover’s performance. We address this by repeating our main evaluation experiment three times and releasing a replication package for practitioners. Second, the operating system environment may affect the evaluation result for the project subjects in SWE-bench lite (e.g., matplotlib). We share the evaluation environment in Section 5 and provide a docker environment for usage by the community. Third, we consider a task instance in SWE-bench lite as resolved if the patch generated by AutoCodeRover passes all the test cases added in the pull request for the issue. However, this patch may be incorrect but overfits the test suite, which is known as a plausible patch in the program repair research community (Gao et al., 2019). We can address this threat by manually verifying whether these patches are semantically equivalent to developers’ patches. Finally, our result of Swe-agent-rep is lower than what is reported by SWE-bench team. This might be due to the different OS environments, different LLM, and parameter setups. To address this threat, we replicated Swe-agent-rep with the default setting in the provided scripts, and we used the same LLM (gpt-4-0125-preview) on Swe-agent-rep as AutoCodeRover, we also released the used replication scripts and log history for Swe-agent-rep.

### v2

@@ Results.
We present a case study on one of the tasks uniquely resolved by ACR-val-sbfl. Figure 8 shows the issue description of django-13964555https://code.djangoproject.com/ticket/32332 and how AutoCodeRover attempts to retrieve code context with and without SBFL. This issue reported a bug when saving django models to a database. A simplified version of the issue description is shown in the first part of Figure 8, in which important parts are highlighted. In this issue, some “hints” (highlighted in red) mentioned are actually distracting factors for a context retrieval agent. For example, the Product and Order classes describe how the bug can be reproduced, and are not classes that cause the bug. Retrieving code context from these two classes (as shown in Figure 8 part 2 “WITHOUT SBFL”) does not yield useful results for forming the context and resolving the issue.
@@
method is actually where the developer chose to fix this bug [^6] https://github.com/django/django/pull/13964/files). Moreover, we observe that the agent does not solely rely on the SBFL results to make API invocations. Instead of searching for methods ranked as top-1 in the SBFL results, the agent searched for the 3rd and 5th ranked methods. These methods are more related to some other hints mentioned in the issue description (highlighted in yellow), and the LLM agent is able to exploit this correlation between the natural language descriptions and the SBFL analysis results. With the correct context collected, AutoCodeRover is then able to draft a patch that resolves this issue. This suggests that an execution-based analysis can complement the agent workflow by revealing information not included in the issue description.
@@ 6.3. RQ3: Challenges on real-life tasks
In this research question, we analyze the task instances in SWE-bench lite that AutoCodeRover failed to resolve, and provide a taxonomy of the issue characteristics to highlight the practical challenges in achieving fully automated software improvement. Our taxonomy consists of challenges in the fault localization stage and patch generation stage. Specifically, for each task, we analyze the best run in the three repetitions, and classify each of the 300 tasks into one of the following:
@@
Figure 9 shows the distribution of the 300 tasks in SWE-bench lite. AutoCodeRover resolves 22.3% of the issues (“Success”), as mentioned in Section 6.1. The fail-to-resolve cases are included in the remaining four categories. In 31.7% of the tasks, AutoCodeRover correctly decided on all patch locations (at the method-level), but did not produce a correct patch (“Wrong patch”). More fine-grained intra-procedural analysis and specification inference techniques can play a significant role in improving these cases, by providing the patch generation agent with more method-level repair guidance. In the other three categories, the fault localization could not pinpoint all the locations to be modified. In 20.7% of the tasks, a patch is generated in the correct file, but at wrong methods / classes in the file (“Wrong location in correct file”). In some of these runs, the developer patch modifies multiple methods, but the generated patch did not modify all of them. In the other categories, a patch could not be generated at the correct file - in 18.7% of the tasks a patch is generated in wrong files, and in 6.7% of the tasks there is no applicable patch (“Wrong file” and “No patch”). We manually inspected some tasks in these two categories, and observed that their issue description mentions few methods / classes / files in the codebase. Instead, some of them contain short examples to reproduce the issue. For these tasks, one possibility is to generate a comprehensive test-suite based on the issue description, and then use execution information of the test-suite (e.g. SBFL) to reveal suspicious program locations. On the other hand, some other tasks do not contain reproducible examples and only consist of natural language descriptions. For these tasks some human involvement might be helpful. The developers could focus on these tasks.
@@ 8. Threats to Validity
Despite the high efficacy AutoCodeRover achieved in SWE-bench, we report a few potential threats to our approach and experiments and discuss how we addressed them. First, LLM is known for its randomness to generate different results in different runs, which may threaten the validity of AutoCodeRover’s performance. We address this by repeating our main evaluation experiment three times and releasing a replication package for practitioners. Second, the operating system environment may affect the evaluation result for the project subjects in SWE-bench (e.g., matplotlib). We share the evaluation environment in Section 5 and provide a docker environment for usage by the community. Third, we consider a task instance in SWE-bench as resolved if the generated patch passes all the test cases added in the pull request for the issue. However, this patch may be overfitting. We addressed this threat by manually verifying whether the patches are semantically equivalent to developers’ patches among the authors. Finally, our result of Swe-agent-rep is lower than reported by SWE-bench team. This might be due to the different OS environments, different LLM, and parameter setups. To address this, we replicated Swe-agent-rep with the default setting in the provided scripts, and we used the same LLM (gpt-4-0125-preview) on Swe-agent-rep as AutoCodeRover. We also ran experiments in the docker provided by SWE-bench team (Table 3); this docker was publicly released on 6 April 2024.